10616202 11.0 cm Units
Answer #1

11 centimeters = 0.11 meters

r=0.11/2 = 0.055m

E=K*2*q / r^2

336 = K*2*q / r^2

q = 336*r^2/2k

q=336*(0.055)^2/(2*9*10^9)

= 5.65 *10^-11 C